OpenStudy (anonymous):

Solve M = 2x + 3y for y

OpenStudy (anonymous):

M-2x = 3y y = (M-2x)/3

OpenStudy (anonymous):

subtract 2x from both sides M-2x=2x-2x+3y 3y=M-2x divide both sides by 3 \[\frac{\cancel{3}y}{\cancel{3}}=\frac{M-2x}{3}\] \[y=\frac{M-2x}{3}\]